250 uL of 1M azide was added to a 25 mL yeast solution.

What was the concentration of azide in the yeast culture? (use proper units)
Answer #1

Here we use the formula:

C1×V1 = C2×V2

C1: starting concentration = 1M

V1: Starting volume = 250 microL = 0.25 mL

C2: Final concentration = ?

V2: Final volume = 25mL + 0.25 mL = 25.25 mL

1 × 0.25 = C2 × 25.25

0.25 / 25.25 = C2

= 0.0099 M

( Rounded off to 0.01 M)